# Human-compatible driving partners through data-regularized self-play reinforcement learning

### v1

@@ Abstract
A central challenge for autonomous vehicles is coordinating with humans. Therefore, incorporating realistic human agents is essential for scalable training and evaluation of autonomous driving systems in simulation. Simulation agents are typically developed by imitating large-scale, high-quality datasets of human driving. However, pure imitation learning agents empirically have high collision rates when executed in a multi-agent closed-loop setting. To build agents that are realistic and effective in closed-loop settings, we propose Human-Regularized PPO (HR-PPO), a multi-agent algorithm where agents are trained through self-play with a small penalty for deviating from a human reference policy. In contrast to prior work, our approach is RL-first and only uses 30 minutes of imperfect human demonstrations. We evaluate agents in a large set of multi-agent traffic scenes. Results show our HR-PPO agents are highly effective in achieving goals, with a success rate of 93%, an off-road rate of 3.5 %, and a collision rate of 3 %. At the same time, the agents drive in a human-like manner, as measured by their similarity to existing human driving logs. We also find that HR-PPO agents show considerable improvements on proxy measures for coordination with human driving, particularly in highly interactive scenarios. We open-source our code and trained agents. Demonstrations of agent behaviors are shared at https://sites.google.com/view/driving-partners.
@@ 1 Introduction
Developing autonomous vehicles (AVs) that are compatible with human driving remains a challenging task, especially given the low margin for error in the real world. Driving simulators offer a cost-effective and safe means to develop and refine autonomous driving systems. The purpose of these simulators is to prepare AVs for real-world deployment, where they must smoothly interact and coordinate with a diverse set of human drivers. Therefore, a crucial aspect of both learning and validation in these simulators involves realistic simulations: the traffic scenarios and other simulation agents with which the controlled AV interacts. To identify where driving policies fall short, it is important to ensure that the simulated traffic conditions and driver agents closely resemble those in the real world (Gulino et al., 2023; Muhammad et al., 2020).
@@ 3.5 Coordinating with human drivers
What makes HR-PPO agents more compatible with the human logs? To find out, we conduct a qualitative analysis. After analyzing the driving behavior of PPO and HR-PPO agents in 50 randomly sampled scenarios, we conclude that the lower collision rates can be attributed to two main factors. First, the HR-PPO agent’s driving style aligns better with human logs, enabling a higher level of anticipation of other agents’ actions. Secondly, HR-PPO agents maintain more distance from other vehicles, which reduces the risk of collisions. A subset of videos are available at https://sites.google.com/view/driving-partners.
@@ 4 Related work
Recent work has shown that augmenting IL with penalties for driving mistakes can create more reliable policies. This has been demonstrated in both closed-loop (Zhang et al., 2023) and open-loop (Lu et al., 2023) settings. Outside of the driving domain, augmenting goal-conditioned multi-agent reinforcement learning (MARL) with a small amount of observational data has been found to improve the likelihood of convergence to the equilibrium under some settings (Lerer & Peysakhovich, 2019; Hu et al., 2022) and has empirically been shown to yield policies more compatible with existing social conventions of the human reference group (Jacob et al., 2022; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR; Bakhtin et al., 2022). Our approach extends these works to the driving setting where it has not yet been investigated in prior work if this type of data-driven regularization is sufficient to enable convergence to a human-compatible policy.
@@ Acknowledgments
This work is funded by the C2SMARTER Center under Grant Number 69A3552348326 from the U.S. Department of Transportation’s University Transportation Centers Program. This work was also supported in part through the NYU IT High-Performance Computing resources, services, and staff expertise. We are thankful to Mert Çelikok, Aditya Makkar, Sam Sokota, Graham Todd, Xieyuan Zhang, and Yutai Zhou, for their feedback on early versions of this draft. We also thank Franklin Yiu, Alex Tang, and Aarav Pandya for visualizations, last-minute debugging, and being all-around helpful.

### v2

@@ Abstract
A central challenge for autonomous vehicles is coordinating with humans. Therefore, incorporating realistic human agents is essential for scalable training and evaluation of autonomous driving systems in simulation. Simulation agents are typically developed by imitating large-scale, high-quality datasets of human driving. However, pure imitation learning agents empirically have high collision rates when executed in a multi-agent closed-loop setting. To build agents that are realistic and effective in closed-loop settings, we propose Human-Regularized PPO (HR-PPO), a multi-agent algorithm where agents are trained through self-play with a small penalty for deviating from a human reference policy. In contrast to prior work, our approach is RL-first and only uses 30 minutes of imperfect human demonstrations. We evaluate agents in a large set of multi-agent traffic scenes. Results show our HR-PPO agents are highly effective in achieving goals, with a success rate of 93%, an off-road rate of 3.5 %, and a collision rate of 3 %. At the same time, the agents drive in a human-like manner, as measured by their similarity to existing human driving logs. We also find that HR-PPO agents show considerable improvements on proxy measures for coordination with human driving, particularly in highly interactive scenarios. We open-source our code and trained agents at https://github.com/Emerge-Lab/nocturne_lab and share demonstrations of agent behaviors at https://sites.google.com/view/driving-partners.
@@ 1 Introduction
Developing autonomous vehicles (AVs) that are compatible with human driving remains a challenging task, especially given the low margin for error in the real world. Driving simulators offer a cost-effective and safe means to develop and refine autonomous driving systems. The purpose of these simulators is to prepare AVs for real-world deployment, where they must smoothly interact and coordinate with a diverse set of human drivers. Therefore, a crucial aspect of both learning and validation in these simulators involves realistic simulations: the traffic scenarios and other simulation agents with which the controlled AV interacts. To identify where driving policies fall short, it is important to ensure that the simulated traffic conditions and driver agents closely resemble those in the real world Gulino et al. (2023); Muhammad et al. (2020).
@@ 3.5 Coordinating with human drivers
What makes HR-PPO agents more compatible with the human logs? To find out, we conduct a qualitative analysis. After analyzing the driving behavior of PPO and HR-PPO agents in 50 randomly sampled scenarios, we conclude that the lower collision rates can be attributed to two main factors. First, the HR-PPO agent’s driving style aligns better with human logs, enabling a higher level of anticipation of other agents’ actions. Secondly, HR-PPO agents maintain more distance from other vehicles, which reduces the risk of collisions. A subset of videos are available at https://sites.google.com/view/driving-partners
@@
3.6 HR-PPO failure cases
We analyzed 100 scenarios each from the train and test datasets in log-replay mode to understand the type of errors HR-PPO agents make. We identify three types of failure modes and describe them below. Example videos for each group are shared on our project page under “HR-PPO failure cases”. Of the 200 sampled scenarios, 6% of the training dataset and 21% of the test dataset has a failure. The failures are broken down as follows:
  1. 1.
    Sharp turns: About 25% of failures result from off-road events due to challenging turns or target positions.
  2. 2.
    Coordination: Approximately 35% of failures are due to collisions resulting from the failure to anticipate human driving log behavior.
  3. 3.
    Setting-related bugs/failures: Around 35% of failures are caused by unreachable target positions or other noise/errors in the dataset.
This indicates that while there is room for improvement in coordination, the majority of failure cases (approximately 60%) are due to dataset bugs or kinematically challenging goal positions or routes. Due to the high collision rate of the BC policy (26-33% combined, of which 12-14% vehicle collisions and 14-19% off-road events), it is difficult to see if it is actually experiencing coordination failures. Therefore, studying the qualitative differences between coordination failures between HR-PPO and BC is left for future work.
@@ 4 Related work
Recent work has shown that augmenting IL with penalties for driving mistakes can create more reliable policies. This has been demonstrated in both closed-loop (Zhang et al., 2023; Wu et al., 2023) and open-loop (Lu et al., 2023) settings. Outside of the driving domain, augmenting goal-conditioned single-agent reinforcement learning has been found to enhance performance in the Arcade Learning Environment (ALE) Hester et al. (2018) and improve the likelihood of convergence to the equilibrium in certain multi-agent learning settings (Lerer & Peysakhovich, 2019; Hu et al., 2022). In multi-agent settings, it has empirically been shown to yield policies more compatible with existing social conventions of the human reference group (Jacob et al., 2022; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR; Bakhtin et al., 2022). Our approach extends these works to the driving setting where it has not yet been investigated in prior work if this type of data-driven regularization is sufficient to enable convergence to a human-compatible policy.
@@ Acknowledgments
This work is funded by the C2SMARTER Center under Grant Number 69A3552348326 from the U.S. Department of Transportation’s University Transportation Centers Program. This work was also supported in part through the NYU IT High-Performance Computing resources, services, and staff expertise. We are thankful to Mert Çelikok, Aditya Makkar, Sam Sokota, Graham Todd, Xieyuan Zhang, and Yutai Zhou, for their feedback on early versions of this draft. We thank Cole Gulino for taking the time to answer our questions. We also thank Franklin Yiu, Alex Tang, and Aarav Pandya for visualizations, last-minute debugging, and being all-around helpful.